Joan can complete a paper route 20 minutes faster than Lisa. If they work together, they can complete the route in 1 hour. How long (in minutes) does it take Joan to complete the rote on her own?
Please show all steps as the 20 minutes may change to 35 mins, 40 mins, etc. every time the homework problem is opened online!

Answer by richwmiller(17219)

1/x+1/(x+35)=1/60
(2 x+35)/(x (x+35)) = 1/60
x = 105